# Towards Efficient Transferable Preemptive Adversarial Defense

### v1

@@ Abstract
Deep learning technology has brought convenience and advanced developments but has become untrustworthy because of its sensitivity to inconspicuous perturbations (i.e., adversarial attacks). Attackers utilize this sensitivity to slightly manipulate transmitted messages. To defend against such attacks, we have devised a strategy for “attacking” the message before it is attacked. This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features. A forward-backward cascade learning algorithm is used to compute protective perturbations, starting with forward propagation optimization to achieve rapid convergence, followed by iterative backward propagation learning to alleviate overfitting. This strategy offers state-of-the-art transferability and protection across various systems. With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses. We have also devised the first to our knowledge effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised. This work provides a new direction to developing active defenses against adversarial attacks.
@@
Transferable preemptive adversarial defense, adaptive reversion attack, active defense, adversarial attack.
@@ I Introduction
Deep learning technologies have been demonstrated to be sensitive to inconspicuous perturbations (i.e., adversarial attacks) [1]. A man-in-the-middle attacker can utilize this weakness to manipulate transmitted messages (e.g., multimedia such as images, audios, and videos). For example, the attacker can change the identity of an enrolled face to that of the attacker when the legal user is uploading it, enabling the attacker to illegally access (i.e., hack into) the system in the future [2, 3]. The attacker can also alter the response by adding slight noise, thereby degrading the user experience (and possibly, for example, causing an accident during autonomous driving) [4]. Defending against such attacks is a critical but unsolved challenge. Passive defenses, on the one hand, generally sacrifice accuracy for security [5, 6, 7, 8, 9]. Active defenses, on the other hand, do not achieve satisfactory efficiency and transferability [10, 11], making them impractical for efficient user interaction. Moreover, their usage raises concerns about user privacy leakage through a backdoor if a provided model is used [12, 13].
Figure 1: A preemptive defense offers transferable protection against a man-in-the-middle attack by preemptively manipulating the message before it is attacked.
We have devised a preemptive defense strategy we call Fast Preemption (illustrated in Figure 1) that addresses these problems. Fast Preemption preemptively “attacks” (manipulates for enhancement) the message being sent immediately upon uploading, protecting it before being attacked. It has three distinct advantages. (i) Like cryptography [14], an active preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation from the receiver, such as decryption. (ii) Fast Preemption allows plug-and-play implementation because it does not require the receiver to use a paired module, alleviating the security and privacy concerns through a backdoor when using another provider’s system. (iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated at the receiver) is breached, all users are affected. In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings. Therefore, Fast Preemption meets five basic requirements:
@@
  • Not require ground truths: for efficient user interaction, it is impractical to label each message manually.
@@
  • Be transferable to other systems: effect of protection must be transferable across various systems, so receiver does not have to use a paired module.
@@
  • We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring the receiver to use a paired module. It is the first to use different models for the classifier and backbone.
@@ III Preemptive Defense Strategy: Fast Preemption
With the preemptive defense strategy, the sent message is preemptively “attacked” (manipulated for enhancement) as soon as it is uploaded, protecting it before being attacked. In this section, we present the framework and algorithm of our Fast Preemption strategy.
@@ III-A1 Adversarial Attack
An adversarial attack leads to misclassification due to subtle perturbations. Given original image $x$ and its ground truth $y$, the adversarial attacker is able to generate an AE by maximizing the cross entropy (CE) of $x$ for label $y$. Therefore, the optimization objective of an adversarial attack for misclassification is
@@ III-A2 Preemptive Defense
In contrast to backward propagation optimization, we newly define “forward propagation optimization” for the preemptive defense strategy. It directly reverses the optimization direction of Equation (2) from maximization to minimization:
@@ III-B Framework
Classifier: The correctness of the input labels determines the correct optimization direction for defense. However, for efficient user interaction, it is impractical to label each message manually. Classifier $f_{c}$ labels these inputs automatically. Therefore, $y$ in Equations (3) and (4) can be replaced by $f_{c}(x)$. The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy. Processing images on the basis of the ground truth can be regarded as using a perfect classifier.
@@ V-C Attacks
Reversion attacks. We ran reversion attacks to evaluate: (i) whether a Fast Preemption defense is irreversible; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective. Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [9]).
@@ VII-A Classifier
In all active defenses, correct labeling guarantees that optimization moves in the correct direction. As demonstrated in Table VI, a classifier with higher accuracy leads to better performance. As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning. Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction. Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.

### v2

@@ Abstract
Deep learning technology has brought convenience and advanced developments but has become untrustworthy because of its sensitivity to inconspicuous perturbations (i.e., adversarial attacks). Attackers may utilize this sensitivity to manipulate predictions. To defend against such attacks, we have devised a proactive strategy for “attacking” the medias before it is attacked by the third party, so that when the protected medias are further attacked, the adversarial perturbations are automatically neutralized. This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features. A forward-backward cascade learning algorithm is used to compute protective perturbations, starting with forward propagation optimization to achieve rapid convergence, followed by iterative backward propagation learning to alleviate overfitting. This strategy offers state-of-the-art transferability and protection across various systems. With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses. We have also devised the first to our knowledge effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised. This work provides a new direction to developing proactive defenses against adversarial attacks. The proposed methodology will be made available on GitHub.
@@
Transferable preemptive adversarial defense, adaptive reversion attack, proactive defense, adversarial attack.
@@ I Introduction
Deep learning technologies have been demonstrated to be sensitive to inconspicuous perturbations (i.e., adversarial attacks) [1]. An adversarial attacker can utilize this weakness to manipulate medias such as images, audios, and videos. For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access (i.e., hack into) the face authentication system [2, 3]. The attacker can also add slight noise on traffic signs, possibly causing an accident during autonomous driving [4]. Defending against such attacks is a critical but unsolved challenge. Passive defenses, on the one hand, generally sacrifice classification accuracy for resistance against threats [5, 6, 7, 8, 9]. Proactive defenses, on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications. Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is not transferable to other systems [10, 11]. Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [12, 13].
Figure 1: A preemptive defense offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Figure 1. Fast Preemption preemptively “attacks” (manipulates for enhancement) the medias from the user side before being attacked by the third party. Specifically, Fast Preemption learns robust features and predicts possible adversarial perturbations, then enhances these robust features and reverses the predicted perturbations for purpose of protection. In this case, when these protected medias are further attacked, these adversarial perturbations will be automatically neutralized. This strategy has three distinct advantages. (i) A proactive preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation for the classification system. (ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system. (iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated at the classification model) is breached, all inputs are affected. In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings. Therefore, Fast Preemption meets four basic requirements:
@@
  • Not require ground truths: It is impractical to label each input manually.
@@
  • Be transferable to other systems: Effect of protection must be transferable across various systems, so users do not have to use a paired module.
@@
  • We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module. It is the first to use different models for the classifier and backbone.
@@ III Preemptive Defense Strategy: Fast Preemption
With the preemptive defense strategy, the image is preemptively “attacked” (manipulated for enhancement), protecting it before being attacked. In this section, we present the framework and algorithm of our Fast Preemption strategy.
@@ III-A1 Adversarial Attack
An adversarial attack leads to misclassification due to subtle perturbations. Given an original image $x$ and its ground truth $y$, the adversarial attacker is able to generate an AE by maximizing the cross entropy (CE) loss of $x$ for label $y$, thereby reducing the model’s confidence in correctly classifying the input as its true label. Therefore, the optimization objective of an adversarial attack for misclassification is
@@ III-A2 Preemptive Defense
In contrast to backward propagation optimization, we newly define “forward propagation optimization” for the preemptive defense strategy. It directly reverses the optimization direction of Equation (2) from maximization to minimization, which in turn enhances the model’s confidence in accurately classifying the input according to its true label:
@@ III-B Framework
Classifier: The correctness of the input labels determines the correct optimization direction for defense. However, for efficient user interaction, it is impractical to label each input manually. Classifier $f_{c}$ labels these inputs automatically. Therefore, $y$ in Equations (3) and (4) can be replaced by $f_{c}(x)$. The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy. Processing images on the basis of the ground truth can be regarded as using a perfect classifier.
@@ V-C Attacks
Reversion attacks. We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective. Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [9]).
@@ VII-A Classifier
In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction. As demonstrated in Table VI, a classifier with higher accuracy leads to better performance. As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning. Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction. Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.